OpenStudy (anonymous):

Solve the given equation by factoring. x^2-14=5x

hero (hero):

x^2 - 5x - 14 = 0 x^2-7x+2x - 14 = 0 x(x-7)+2(x-7)=0 (x-7)(x+2)=0 x = -2,7